OpenStudy (anonymous):

The ratio of boys to girls in a certain classroom was 2:3 . If boys represented 5 more than one third of class. how many people were in the classroom? a.15 b.25 c.30 d.45

OpenStudy (anonymous):

Let x = boys let y = girls let class = x+ y x/y = 2/3 y = 3/2 (x) boys = 1/3 class + 5 x = (1/3) (x+y) + 5 Now substitute y in for x x = 1/3 (x + 3/2 x) + 5 x = 1/3 x +1/2 x + 5 x = 2/6 x + 3/6 x + 5 x = 5/6 x + 5 1/6 x = 5 x = 30 y = 3/2 (x) y = 3/2 (30) y = 45 class = x+y = 30 +45 = 75 students.

OpenStudy (anonymous):

Oh.. Thanks .. That was very Helpful. ^^ .. I guess the question I have is the wrong one. Your answer seems totally logical.

OpenStudy (anonymous):

Hope it helps! Sorry I tried to get the answers on your multiple choice but maybe they typed the wrong question? o.o Sorry...

OpenStudy (anonymous):

No. Its alright. You are not to blame,friend. I was confused as well. But since you got the same answer.. I guess Mr.Question is the one to blame.